OpenStudy (hellomynameiskam):

What is 4 3/8 - 1/4? A. 4 1/8 B. 4 1/3 C. 4 1/4 D. 4 1/2

OpenStudy (shishi990):

a

OpenStudy (anonymous):

well first subtract the whole numbers witch means 4-1=3, then you make the fractions have a comon denominator. We can use 8

OpenStudy (anonymous):

you need to turn them into improper fractions

OpenStudy (anonymous):

4 3/8=35/8 and 1/4 you can leave that one alone

OpenStudy (anonymous):

you would be subtracting 4 3/8-2/8 and that gives you your answer

OpenStudy (anonymous):

or you could do that :D either way(s) work

OpenStudy (anonymous):

so 35/8-1/4 you need to change to a common denominator so 1/4=2/8

OpenStudy (anonymous):

exactly and then simplify

OpenStudy (anonymous):

then subtract 35-2=33 put it over 8 and you have 33/8 simplify and you have 4 1/8

OpenStudy (anonymous):

yep that is your answer